calculate pH and pOH of a hydrochloric acid solution prepared by dissolving 30.5kg of hygrogen chloride gas in 806L of water at SATP
1
Expert's answer
2018-01-30T14:12:08-0500
n(HCl)=(m(HCl))/(M(HCl))=(30.5 kg)/(36.5 g/mol)=(30.5×〖10〗^3 g)/(36.5 g/mol)=835.6 mol
[HCl]=(n(HCl))/(V(solution))=(835.6 mol)/(806 L)=1.04 M
pH=-lg⁡[H^+ ]=-lg⁡[HCl]=-0.0170
pOH=14-pH=14+0.0170=14.0170
Answer:
pH = - 0.0170
pOH = 14.0170